OpenStudy (anonymous):

Which answer choice is equivalent to the expression below? 3x2y4/2xy The 2 and 4 in the first part of problem(not the 2 from 2xy) are exponents. A.3xy3/2 ( The second 3 is exponent ) B.3x3y5/2 (The second 3 is exponent, and the 5) C.3/2xy3 ( The 3 in 2xy3 is exponent) D.3xy2/2 ( The 2 in 3xy2 is an exponent.)

OpenStudy (kirbykirby):

A. Just use subtraction on the exponents to know how many to get rid of. so (y^4)/y = y^3 since you do 4-1 on the exponent. , same for (x^2)/x = x , so you're left with answer A